OpenStudy (anonymous):

A family refers to a group of two or more people related by birth, marriage, or adoption who reside together. In 2000, in Country A, the average family net worth was $310,000, and there were about 8.3 x 10 (7 power) families. Calculate the total family net worth in Country A in 2000. I am stuck here: 3.1x 10(fifth power) x 8.3 x 10 (7th power) = 3.1 x 8.3 = 25.73 = 10(fifth power) x 10 (7th power) = 10 (12th power) =

OpenStudy (mayankdevnani):

\[\huge 3,10,000=3.1 \times 10^5\] and \[\huge 8.3 \times 10^7\] just multiply them

OpenStudy (mayankdevnani):

\[\huge 3.1 \times 10^5 \times 8.3 \times 10^7\] \[\huge 3.1 \times 8.3 \times 10^5 \times 10^7\]

OpenStudy (mayankdevnani):

can you solve it?? @mister88

OpenStudy (anonymous):

That is what I have done so far. 25.73 10(12 power) 25.73 x 10 (12 power) 2,573 plus the 12?

OpenStudy (mayankdevnani):

so answer will look like this : \[\huge 25.73 \times 10^{12}\] it is also written as : \[\huge 2.573 \times 10^{12-1}=2.573 \times 10^{11}\]

OpenStudy (mayankdevnani):

hope you understand . @mister88

OpenStudy (anonymous):

I wasn't sure about adding or subtracting the one?

OpenStudy (mayankdevnani):

no, when we move the decimal point to left by one place, the power of 10 will decrease by 1

OpenStudy (mayankdevnani):

understood or not? @mister88

OpenStudy (anonymous):

Okay that is the part that I forgot about....Thanks

OpenStudy (mayankdevnani):

welcome :)